# Classical and quantum chaotic synchronization in coupled dissipative time crystals

### v2

@@ 5 Conclusions
In summary, this model displays classical chaotic synchronization and an analogous quantum phenomenon, which we can call quantum chaotic synchronization. Interestingly, the crossover point between the staggered and the uniform magnetization in the mean-field classical chaotic synchronization is consistent with the one in the quantum regime, confirming quantum-classical correspondence. This is an interesting finding because the limits $S\to\infty$ and $t\to\infty$ do not commute in this quantum dynamics. In the mean-field case, the $S\to\infty$ limit is taken first, resulting in persistent oscillations that can be periodic or aperiodic [39]. In the quantum case, the $t\to\infty$ limit comes first and the system always relaxes to a NESS. Remarkably, the NESS shares qualitative and quantitative features with the mean-field state, despite their fundamental differences and the presence of entanglement only in the quantum case. The noncommutativity of these limits explains the discrepancies in parameter regimes where classical and quantum chaos appear.
@@
We acknowledge P. Solanki for fruitful discussions. G. P. and A. R. acknowledge financial support from PNRR MUR Project No. PE0000023-NQSTI. We acknowledge computational resources from the CINECA award under the ISCRA initiative and from MUR, PON “Ricerca e Innovazione 2014-2020,” under Grant No. PIR01_\_00011 - (I.Bi.S.Co.). This work was supported by PNRR MUR Project No. PE0000023-NQSTI, by the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 101017733, by MUR Project No. CN_\_00000013-ICSC (P. L.), and by the QuantERA II Programme STAQS project, which received funding from the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 101017733 (P. L.).

### v4

@@ 5 Conclusions
In summary, this model displays classical chaotic synchronization and an analogous quantum phenomenon, which we can call quantum chaotic synchronization. Interestingly, the classical crossover point between the staggered and the uniform magnetization – where chaotic synchronization sets in – does not in general coincide with the corresponding quantum one. This is related to the fact that the limits $S\to\infty$ and $t\to\infty$ do not commute with each other. In the mean-field case, the $S\to\infty$ limit is taken first, resulting in persistent oscillations that can be periodic or aperiodic [39]. In the quantum case, the $t\to\infty$ limit comes first and the system always relaxes to a NESS. Remarkably, the NESS and the mean-field dynamics share a qualitatively similar chaotic synchronization behavior, despite their fundamental differences and the presence of entanglement only in the quantum case. The noncommutativity of these limits furthermore explains the discrepancies in parameter regimes where classical and quantum chaos appear.
@@
We acknowledge P. Solanki for fruitful discussions. G. P. and A. R. acknowledge financial support from PNRR MUR Project No. PE0000023-NQSTI. We acknowledge computational resources from the CINECA award under the ISCRA initiative and from MUR, PON “Ricerca e Innovazione 2014-2020,” under Grant No. PIR01_\_00011 - (I.Bi.S.Co.). This work was further supported by the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 101017733, by MUR Project No. CN_\_00000013-ICSC (P. L.), and by the QuantERA II Programme STAQS project, which received funding from the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 101017733 (P. L.). E. P. acknowledges financial support from the Erasmus+ AURORA ALLIANCE fellowship program. E. P. acknowledges support of the Czech Science Foundation (project 25-17472S). E. P. acknowledges project IGA PrF-2025-010.